OpenStudy (maheshmeghwal9):

Prove it plz :) \[\LARGE{1^2+2^2+3^2+.....+n^2> \frac{n^3}{3}}\]

OpenStudy (maheshmeghwal9):

@UnkleRhaukus @hartnn @Hero @satellite73 @amistre64 @Callisto @experimentX Please help:D

OpenStudy (maheshmeghwal9):

@lalaly too:)

hartnn (hartnn):

do you know standard formula for 1^2+2^2+3^2+.... n^2 =... ?

OpenStudy (ash2326):

Sum of \(\large 1^2+2^2+3^2...n^2=\frac{(n)(n+1)(2n+1)}{6}\) If you expand this, you can prove it easily. Can you do that @maheshmeghwal9??

OpenStudy (maheshmeghwal9):

no :(

OpenStudy (maheshmeghwal9):

i mean how to expand?

OpenStudy (ash2326):

Multiply the terms :) \[(n)(n+1)(2n+1)\]

OpenStudy (maheshmeghwal9):

ok then i gt this \[3n^2+2n^3+n\]

OpenStudy (ash2326):

Divide this by 6,

OpenStudy (maheshmeghwal9):

ok then here it is \[\frac{n^2}{2}+\frac{n^3}{3}+\frac{n}{6}.\]

OpenStudy (ash2326):

Obviously for positive n \[\large \frac{n^3}{3}+\frac{n^2}{2}+\frac{n}{6} > \frac {n^3}{3} \]

OpenStudy (maheshmeghwal9):

yeah thanx:)

Parth (parthkohli):

How do you conclude that it is “obvious” for a positive \(n\)? @ash2326

OpenStudy (maheshmeghwal9):

@abhyudaysingh12 got it or not?

OpenStudy (maheshmeghwal9):

for every 'n' it is true actually:)

Parth (parthkohli):

How would you conclude that?

hartnn (hartnn):

\(\large \frac{n^2}{2}+\frac{n}{6} >0\) for n>0

hartnn (hartnn):

^ that way

hartnn (hartnn):

then add n^3/3 on both sides.

Parth (parthkohli):

\[n^2 + n > 0 \iff n(n + 1)>0 \iff n +1>0\iff n>-1\]

OpenStudy (maheshmeghwal9):

BUT question says that n>0 so n>-1 is ignored 1,2,.........n see @ParthKohli :)

Parth (parthkohli):

But you asserted that it's true for ALL \(n\).

OpenStudy (maheshmeghwal9):

sorry for that statement but that i gt in haste;)

Parth (parthkohli):

lol okay

OpenStudy (maheshmeghwal9):

:D good job;)

Parth (parthkohli):

?

OpenStudy (maheshmeghwal9):

to discuss something is good job:)

Parth (parthkohli):

?

OpenStudy (maheshmeghwal9):

?

OpenStudy (experimentx):

there are couple of ways you can do it ... few of them are above. Apart from that, you can also try induction.

OpenStudy (unklerhaukus):